OpenStudy (mathhelp346):

Write the equation in standard form of the line that is parallel to -6x-3y=-2 and contains (8,-2)

OpenStudy (anonymous):

A parallel line has the same slope, a perpendicular line has a negative inverse for the slope.

OpenStudy (mathhelp346):

Ok, so the slope is -2

OpenStudy (anonymous):

2x+1y=14? Just guessing!

OpenStudy (anonymous):

Solve for y to determine what the slope of this equation is.

OpenStudy (mathhelp346):

Yes the slope is -2

OpenStudy (anonymous):

Sorry, I'm a bit slow today. Yes, you are correct.

OpenStudy (anonymous):

Use the point-slope formula to write the new equation.

OpenStudy (mathhelp346):

y=-2x+2/3 Mia how did you get that? I graphed it and it was correct

OpenStudy (anonymous):

\(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

Well you said it had to be in standard form so I just reorded it to fit standard form (Ax+By=C).

OpenStudy (mathhelp346):

How did you reorder it?

OpenStudy (anonymous):

Well I got y=-2x+14. Move the -2x by adding it to both sides getting 2x+y=14.(:

OpenStudy (mathhelp346):

Ohh I see thank you

OpenStudy (anonymous):

You're welcome(: